# Accretion onto supermassive and intermediate mass black holes in cosmological simulations

### v1

@@
Accretion onto supermassive and intermediate mass black holes in cosmological simulations
Accretion is the dominant contribution to the cosmic massive black hole density in the Universe today. Yet, modelling it in cosmological simulations is challenging due to the dynamic range involved, as well as the theoretical uncertainties of the underlying mechanisms driving accretion from galactic to black hole horizon scales. We present a simple, flexible parametrization for gas inflows onto massive black holes in order to manage this uncertainty in large-volume cosmological simulations. This is done as part of the “Learning the Universe” collaboration, which aims to jointly infer the initial conditions and physical processes governing the evolution of the Universe using a Bayesian forward-modelling approach. To allow such a forward-modelling, we update the prescription for accretion with a two-parameter free-fall based inflow estimate that allows for a radius-dependent inflow rate and add a simple model for unresolved accretion disks. We use uniform resolution cosmological hydrodynamical simulations and the IllustrisTNG framework to study the massive black hole population and its dependence on the introduced model parameters. Once the parameters of the accretion formula are chosen to result in a roughly similar redshift zero black hole mass density, the differences caused by the details in the accretion formula are moderate in the supermassive black hole regime, indicating that it is difficult to distinguish between accretion mechanisms based on luminous active galactic nuclei powered by supermassive black holes. Applying the same models to intermediate mass black holes at high redshift, however, reveals significantly different accretion rates in high redshift, moderate luminosity active galactic nuclei and different frequencies and mass distributions of intermediate mass black hole mergers for the same black hole formation model. This difference in early growth history will likely also lead to an accretion model dependent supermassive black hole population in the case of light massive black hole formation scenarios.
@@
Massive black holes (MBHs) are among the most poorly understood objects in extragalactic astrophysics. While their existence can be established using stellar dynamics in the galactic centre (Ghez et al. 2008; Genzel et al. 2010) or emission of the inner accretion disk (Event Horizon Telescope Collaboration et al. 2019), their evolution is relatively poorly constrained (Greene et al. 2020; Inayoshi et al. 2020; Harikane et al. 2023; Taylor et al. 2024; Matthee et al. 2024). Studies of luminous active galactic nuclei (AGNs) show that most of their overall mass growth is dominated by gas accretion in bright quasi-stellar object phases (Soltan 1982; Yu & Tremaine 2002), with the hard X-ray background dominated by the cosmic black hole accretion rate density (Ueda et al. 2003; Shankar et al. 2009). Comparing its redshift dependence to the star formation rate density (Merloni et al. 2004) hints towards a co-evolution picture of supermassive black holes and their host galaxies, with the most massive MBHs assembling earlier than the lower mass ones (Shankar et al. 2004).
@@
Mergers with other MBHs are a natural consequence of hierarchical structure formation and represent a second growth channel. While overall subdominant (Ni et al. 2022), this channel can contribute substantially to the mass growth of MBHs in specific evolutionary stages, assuming the inspiral time of a MBH binary from kpc scales is sufficiently rapid: since the host galaxy merger rate is dictated by cosmological structure formation, this primarily happens in phases where gas accretion is suppressed such as during early growth (Bhowmick et al. 2024b) and late time evolution in massive galaxies and galaxy clusters (Weinberger et al. 2018).
@@
The masses of MBHs show correlations with the stellar content of their host galaxies (Magorrian et al. 1998), nuclear kinematics (Ferrarese & Merritt 2000; Gebhardt et al. 2000; McConnell & Ma 2013) and other galactic properties (e.g. Graham & Driver 2007; Martín-Navarro et al. 2018). How to precisely interpret these correlations is still debated (Kormendy & Ho 2013), with galaxy mergers generally being seen to play a central role in them: on the one hand, they cause MBH mergers, which in turn narrow the distribution of massive black hole to stellar mass ratios (Peng 2007; Jahnke & Macciò 2011; Graham & Sahu 2023). On the other hand, they trigger central starbursts and possibly quasar activity (Di Matteo et al. 2005; Hopkins et al. 2006, 2008), manifesting in correlations between luminous AGN and mergers (Ellison et al. 2019; Marian et al. 2020). Yet, even when taking the cosmological formation history into account, the interpretation of these correlations still depends on the assumed gas accretion rate from galactic to accretion disk scales, most notably its MBH mass dependence (Anglés-Alcázar et al. 2013).
@@
In the early universe, high redshift quasars powered by rapidly accreting supermassive black holes are detected up to and exceeding (see Inayoshi et al. 2020; Fan et al. 2023, and references therein). These objects represent the most extreme formation and growth scenarios of MBHs, with hints of a significantly more abundant population of lower mass, high redshift AGN recently being found due to better observational capabilities (Maiolino et al. 2023; Adamo et al. 2024; Greene et al. 2024; Kokorev et al. 2024; Kocevski et al. 2024). While the mass estimates of the MBHs powering these sources are highly uncertain, the existence of these luminous AGNs seems to be a challenge to commonly assumed evolutionary pathways (Akins et al. 2024; Durodola et al. 2024; Kovács et al. 2024). This concerns in particular the MBH-galaxy scaling relations of these objects, indicating that these MBHs are substantially over-massive, even when taking the substantial uncertainties in MBH and stellar mass measurements into account (Pacucci et al. 2023; Natarajan et al. 2024).
@@
At the opposite extreme are MBHs that never undergo any substantial growth period and stay at relatively similar masses until (see Greene et al. 2020, for a review). These intermediate mass black holes (IMBHs) should be found in dwarf galaxies or globular clusters (Mezcua 2017) with unambiguous evidence being challenging to obtain (see however Häberle et al. 2024). From a MBH evolution perspective, these “leftover” IMBHs are the slowest growing and thus cleanest remnants of the formation process of MBHs in the high redshift Universe, thus a promising avenue to distinguish between possible formation scenarios (see Volonteri et al. 2021, for a review).
@@
Understanding accretion onto MBHs from a theoretical perspective is challenging not only due to the dynamic range from galactic to black hole scales, but also due to a variety of other physical processes that simultaneously happen in galactic centres. Yet, substantial progress has been made in the past few years simulating accretion flows from galactic to accretion disk scales, in large parts driven by advances in adaptive refinement (Curtis & Sijacki 2015; Anglés-Alcázar et al. 2021) and improved algorithms to cover the necessary dynamic range. Studies of quasars (Curtis & Sijacki 2016; Hopkins et al. 2024a, b; Lupi et al. 2024) as well as accretion flows onto elliptical galaxies (Guo et al. 2023, 2024), galaxy groups and clusters (Gaspari et al. 2017), and Milky-Way mass galaxies (Emsellem et al. 2015; Ressler et al. 2018) have all demonstrated that it is possible to study specific environments of MBHs, gaining substantial insight into the nature of accretion flows. Yet, gas inflows in these different environments seem to be dominated by different aspects (cooling time, gravitational torques, gas supply, magnetic fields, etc.) and a unified picture is not yet emerging.
@@
In cosmological simulations of galaxy formation, MBHs have become an essential ingredient (see Di Matteo et al. 2023, and references therein). The AGN feedback invoked in the simulation models proves critical for reproducing the high mass end of the galaxy population (Sijacki et al. 2007; Weinberger et al. 2017), galaxy cluster properties (Barnes et al. 2018) and even affects cosmological structure formation itself (Springel et al. 2018). Solving for the cosmological evolution of halos, their host galaxies with gas cooling, star formation and gas inflows onto the central MBHs over cosmic time yields a wealth of predictions about massive black holes and their host galaxies which can be compared to observations (Habouzit et al. 2019; Terrazas et al. 2020; Li et al. 2020), making it a powerful predictive tool, provided the employed model is realistic and accurate.
@@
One notable incompleteness of many current state-of-the-art galaxy formation models are IMBHs and the ability to represent formation channels that result in lower-mass “seed” black holes that could range from M⊙. This is partially due to computational limitations, where the mass of a resolution element is often larger than the respective formation mass. However, the MBH model might itself be an obstacle due to the employment of the Bondi-Hoyle formula (Bondi & Hoyle 1944; Bondi 1952) as the accretion estimate (see however Davé et al. 2019, who use a different accretion model). The quadratic scaling of the accretion rate with black hole mass leads to a suppression of accretion for low-mass IMBHs, especially when employed in combination with an effective equation of state model for the interstellar medium that provides a floor to the sound speed (Booth & Schaye 2009). This leads to a growth history that is too delayed to explain high redshift luminous AGN unless massive seeds are invoked. While IMBHs can still grow significantly via mergers in permissive seeding models that predict frequent seed formation (Bhowmick et al. 2024a), an alternative solution to this problem cold be that the Bondi-Hoyle formula is simply inadequate for this early growth regime. Since neither our theoretical understanding nor our observational capabilities are able to rule out either scenario directly, there is a need to model both in cosmological simulations to explore indirect effects.
@@
This work is part of the “Learning the Universe” collaboration111https://www.learning-the-universe.org dedicated to understanding the extragalactic Universe by jointly inferring the initial conditions and physical laws that govern its subsequent evolution. As part of the effort to build the predictive galaxy formation model needed for this endeavour, we present a new, flexible parametrization for gas inflows and MBH growth via accretion in large volume cosmological simulations. The model starts out by measuring the gas properties at a fixed, galactic scale, assumes a scaling of the mass inflow rate with radius, and feeds a simple prescription for accretion disk evolution, which ultimately grows the MBH.
@@ 2.1 Initial conditions
We run uniform resolution cosmological simulations starting at , created adopting the Planck Collaboration et al. (2016) cold dark matter cosmological parameters (, , , km s-1 Mpc km s-1 Mpc-1, and ) and using the same linear theory power spectrum as the IllustrisTNG simulations. The initial displacement is calculated using the NgenIC code in the version originally developed for Gadget4 (Springel et al. 2021) using second-order Lagrangian perturbation theory and applying a variance suppression technique (Angulo & Pontzen 2016) to enhance the representativeness of the simulated volume. Gas is initialized from the same density and velocity field, splitting the matter into an equal number of gas cells and dark matter particles and distributing the mass according to the cosmic baryon and dark matter fractions, respectively. A uniform magnetic field is initialized with comoving Gauss.
In the following, we present cosmological simulations with two different volumes: to understand the overall evolution over cosmic time, we evolve boxes of sidelength Mpc (L50) to redshift probing structure formation of relatively massive halos, but at comparably low mass resolution. To study the high-redshift, early growth of intermediate mass black holes (IMBHs), we evolve smaller volumes with sidelength of Mpc (L12.5) to redshift . These simulations adopt lower MBH seed masses and are used to study the IMBH growth at high redshifts. Throughout this paper, we use MBH as the general term encompassing both IMBHs with mass M⊙ and SMBHs with M⊙. All simulations are run with , and initial resolution elements, covering overall a dynamic range of in mass resolution and an identical mass resolution of the lowest resolution L12.5 and the highest resolution L50 simulation. Unless stated explicitly, we show the highest resolution version of the respective simulation box.
@@ 2.2 Reference: the IllustrisTNG model
The simulations are carried out using the Arepo code (Springel 2010; Pakmor et al. 2011, 2016; Weinberger et al. 2020) solving the equations of magnetohydrodynamics using the finite-volume technique on a moving mesh and the Poisson equation using a tree-particle-mesh method. An 8-wave Powell cleaning scheme is applied to ensure approximately divergence-free magnetic fields (Pakmor & Springel 2013).
@@
Our main aim is to develop and test a new model for accretion onto MBHs in cosmological environments. We use the IllustrisTNG galaxy formation model as a well-studied framework with comprehensive analysis of the behaviour of its simulated MBH population in the literature (e.g. Bhowmick et al. 2020; Terrazas et al. 2020; Li et al. 2020; Truong et al. 2021; Habouzit et al. 2021, 2022a, 2022b, among many others). In the following, we briefly summarize the IllustrisTNG model. For a more detailed description, we refer the reader to the IllustrisTNG method papers (Weinberger et al. 2017; Pillepich et al. 2018).
@@
We include radiative cooling down to K using primoridal (Katz et al. 1996) and metal-line cooling (Wiersma et al. 2009), taking into account a time dependent spatially homogeneous UV background (Faucher-Giguère et al. 2009)222Note that revised backgrounds exist (Puchwein et al. 2019; Faucher-Giguère 2020). However, to be consistent with IllustrisTNG we use the original model. and adding corrections for self-shielding (Rahmati et al. 2013). Additionally, the radiation field of nearby AGN is taken into account. The implementation details are described in Vogelsberger et al. (2013).
@@
Star formation is modelled following Springel & Hernquist (2003); Vogelsberger et al. (2013), with dense gas above a threshold density of cm-3 being star forming and on an effective equation of state (or in a hotter state), thereby modelling the unresolved interstellar medium. Star particles form in a stochastic manner from star forming gas and are evolved assuming a single stellar population with Chabrier (2003) initial mass function, taking into account chemical enrichment and mass return from asymptotic giant branch stars, core-collapse and type Ia supernovae (with the TNG model updates as described in Pillepich et al. 2018). Stellar feedback is parameterized via massive, enriched galactic wind particles that are temporarily decoupled from the gas dynamics (Springel & Hernquist 2003; Vogelsberger et al. 2013) using the scaling of the IllustrisTNG simulation (Pillepich et al. 2018).
@@
The MBH modelling follows a relatively simple, numerically robust approach for seeding, dynamics and mergers (Springel et al. 2005; Sijacki et al. 2007; Di Matteo et al. 2008). Specifically, in every halo with a mass exceeding M⊙ that does not yet contain one, a MBH is seeded converting the most bound gas cell into the black hole particle with seed mass M⊙. This black hole particle is kept at the potential minimum of the halo throughout the simulation, with MBH mergers being modelled instantaneously without explicitly following the dynamics of the black holes.
@@
Gas accretion onto MBHs is modelled using the Eddington-limited Bondi formula applied to a kernel-averaged estimate of the gas density and sound speed in the cells surrounding the black hole particle (average denoted by ).
@@
where is the gravitational constant, the massive black hole mass, the proton mass, the radiative efficiency, the Thompson cross-section and the speed of light. The corresponding black hole mass growth is
@@
The averaging scale is adjusted on-the-fly at every timestep using a target weighted number of cells
@@
with a given tolerance333Note that we keep this number fixed with resolution, which is a deviation from the IllustrisTNG simulation, where it is scaled by a factor of for each resolution step (factor of in target mass).. denotes the distance of the cell mesh-generating point and the MBH position and the target gas mass (i.e., the gas mass resolution). The weighting kernel is defined as
@@
Feedback from AGNs is modelled in two modes (Weinberger et al. 2017). Highly accreting MBHs relative to the Eddington limit inject thermal energy continously into their surroundings (using the same kernel)
For low accretion rate MBHs, a kinetic kick is injected in a pulsed fashion whenever enough energy is available.
Note that both feedback channels are mutually exclusive, with the threshold being

### v2

# Accretion onto supermassive and intermediate-mass black holes in cosmological simulations

@@
Accretion onto supermassive and intermediate-mass black holes
in cosmological simulations
Accretion is the dominant contribution to the cosmic massive black hole (MBH) density in the Universe today. However, modelling accretion in cosmological simulations is challenging due to the dynamic range involved, as well as the theoretical uncertainties of the underlying mechanisms driving accretion from galactic to black hole horizon scales. We present a simple, flexible parametrisation for gas inflows onto MBHs aimed at managing this uncertainty in the context of large-volume cosmological simulations. This study has been carried out as part of the ‘Learning the Universe’ collaboration, with an aim to jointly infer the initial conditions and physical processes governing the evolution of the Universe, using a Bayesian forward-modelling approach. To allow for this forward-modelling approach, we updated the prescription for accretion with a two-parameter free-fall based inflow estimate that allows for a radius-dependent inflow rate and added a simple model for unresolved accretion disks. We used uniform resolution cosmological hydrodynamical simulations and the IllustrisTNG framework to study the MBH population and its dependence on the introduced model parameters. Once the parameters of the accretion formula were chosen, aimed at achieving a zero black hole mass density (BHMD) at a roughly similar redshift, the differences caused by details in the accretion formula are moderate in the supermassive black hole (SMBH) regime, indicating that it is difficult to distinguish between accretion mechanisms based on luminous active galactic nuclei (AGNs) powered by SMBHs. Applying the same models to intermediate-mass black holes (IMBHs) at high redshifts, however, reveals significantly different accretion rates in high-redshift, moderate-luminosity AGNs, as well as different frequencies and mass distributions of IMBH mergers for the same black hole formation model. This difference in the early growth history will also likely lead to an accretion model-dependent SMBH population in low-mass black hole formation scenarios.
@@
Massive black holes (MBHs) are among the most poorly understood objects in extragalactic astrophysics. While their existence can be established using stellar dynamics in the galactic centre (Ghez et al. 2008; Genzel et al. 2010) or emission of the inner accretion disk (Event Horizon Telescope Collaboration et al. 2019), their evolution is relatively poorly constrained (Greene et al. 2020; Inayoshi et al. 2020; Harikane et al. 2023; Taylor et al. 2024; Matthee et al. 2024). Studies of luminous active galactic nuclei (AGNs) show that most of their overall mass growth is dominated by gas accretion in bright quasi-stellar object phases (Soltan 1982; Yu & Tremaine 2002), with the hard X-ray background dominated by the cosmic black hole accretion rate density (BHARD; Ueda et al. 2003; Shankar et al. 2009). Comparing its redshift dependence to the star formation rate density (Merloni et al. 2004) hints at a likely co-evolution scenario for supermassive black holes (SMBHs) and their host galaxies, with the most massive MBHs assembling earlier than the lower mass ones (Shankar et al. 2004).
@@
Mergers with other MBHs are a natural consequence of hierarchical structure formation and represent a second growth channel. While overall subdominant (Ni et al. 2022), this channel can contribute substantially to the mass growth of MBHs in specific evolutionary stages, assuming the inspiral time of a MBH binary from kpc scales is sufficiently rapid: since the host galaxy merger rate is dictated by cosmological structure formation, this primarily happens in phases where gas accretion is suppressed such as during early growth (Bhowmick et al. 2024b) and late-time evolution in massive galaxies and galaxy clusters (Weinberger et al. 2018).
@@
The masses of MBHs show correlations with the stellar content of their host galaxies (Magorrian et al. 1998), nuclear kinematics (Ferrarese & Merritt 2000; Gebhardt et al. 2000; McConnell & Ma 2013), and other galactic properties (e.g. Graham & Driver 2007; Mart´ın-Navarro et al. 2018). The question of how we can interpret these correlations is still a matter of debate (Kormendy & Ho 2013), with galaxy mergers generally being seen to play a central role in them. On the one hand, they cause MBH mergers, which, in turn, narrow the distribution of MBHs to stellar mass ratios (Peng 2007; Jahnke & Macciò 2011; Graham & Sahu 2023). On the other hand, they trigger central starbursts and (possibly) quasar activity (Di Matteo et al. 2005; Hopkins et al. 2006, 2008), manifested as correlations between luminous AGNs and mergers (Ellison et al. 2019; Marian et al. 2020). Yet, even when taking the cosmological formation history into account, the interpretation of these correlations still depends on the assumed gas accretion rate from galactic to accretion disk scales, most notably its MBH mass dependence (Anglés-Alcázar et al. 2013).
@@
In the early Universe, high redshift quasars powered by rapidly accreting SMBHs are detected up to and exceeding (see Inayoshi et al. 2020; Fan et al. 2023, and references therein). These objects represent the most extreme formation and growth scenarios of MBHs, with hints of a significantly more abundant population of lower mass, high redshift AGN recently being found due to better observational capabilities (Maiolino et al. 2024; Adamo et al. 2024; Greene et al. 2024; Kokorev et al. 2024; Kocevski et al. 2024). While the mass estimates of the MBHs powering these sources are highly uncertain, the existence of these luminous AGNs seems to be a challenge to commonly assumed evolutionary pathways (Akins et al. 2024; Durodola et al. 2024; Kovács et al. 2024). This concerns in particular the MBH-galaxy scaling relations of these objects, indicating that these MBHs are substantially over-massive, even when taking the substantial uncertainties in MBH and stellar mass measurements into account (Pacucci et al. 2023; Natarajan et al. 2024).
@@
At the opposite extreme, we have MBHs that have never undergone any substantial growth period and remain at relatively similar masses until (see Greene et al. 2020, for a review). These intermediate-mass black holes (IMBHs) are expected to be found in dwarf galaxies or globular clusters (Mezcua 2017); however, unambiguous evidence is still lacking (see, however, Häberle et al. 2024). From a MBH evolution perspective, these ‘leftover’ IMBHs are the slowest growing and thus cleanest remnants of the formation process of MBHs in the high redshift Universe, thus a promising avenue to distinguish between possible formation scenarios (see Volonteri et al. 2021, for a review). Throughout this paper, we use MBH as the general term encompassing both IMBHs with mass M⊙ and SMBHs with M⊙.
@@
Understanding accretion onto MBHs from a theoretical perspective is challenging not only due to the dynamic range from galactic to black hole (BH) scales, but also due to a variety of other physical processes that simultaneously happen in galactic centres. However, substantial progress has been made in the past few years simulating accretion flows from galactic to accretion disk scales, in large parts driven by advances in adaptive refinement (Curtis & Sijacki 2015; Anglés-Alcázar et al. 2021) and improved algorithms to cover the necessary dynamic range. Studies of quasars (Curtis & Sijacki 2016; Hopkins et al. 2024a, b; Lupi et al. 2024) as well as accretion flows onto elliptical galaxies (Guo et al. 2023, 2024), galaxy groups and clusters (Gaspari et al. 2017), and Milky-Way mass galaxies (Emsellem et al. 2015; Ressler et al. 2018) have all demonstrated that it is possible to study specific environments of MBHs. These approaches have provided substantial insights into the nature of accretion flows. Still, gas inflows in these different environments seem to be dominated by different aspects (cooling time, gravitational torques, gas supply, magnetic fields, etc.) and a unified picture has not yet emerged.
@@
In cosmological simulations of galaxy formation, MBHs have become an essential ingredient (see Di Matteo et al. 2023, and references therein). The AGN feedback invoked in the simulation models proves critical for reproducing the high mass end of the galaxy population (Sijacki et al. 2007; Weinberger et al. 2017) and determining galaxy cluster properties (Barnes et al. 2018). It even affects cosmological structure formation per se (Springel et al. 2018). Solving for the cosmological evolution of halos, their host galaxies with gas cooling, star formation, and gas inflows onto the central MBHs over cosmic time yields a wealth of predictions about MBHs and their host galaxies which can be compared to observations (Habouzit et al. 2019; Terrazas et al. 2020; Li et al. 2020). This makes MBHs a powerful predictive tool, provided the employed model is realistic and accurate.
@@
One notable incompleteness of many current state-of-the-art galaxy formation models are IMBHs and the ability to represent formation channels that result in lower-mass ‘seed’ BHs that could range from M⊙. This is partially due to computational limitations, where the mass of a resolution element is often larger than the respective formation mass. However, the MBH model might itself be an obstacle due to the employment of the Bondi-Hoyle formula (Bondi & Hoyle 1944; Bondi 1952) as the accretion estimate (see however Davé et al. 2019, who use a different accretion model). The quadratic scaling of the accretion rate with BH mass leads to a suppression of accretion for low-mass IMBHs, especially when employed in combination with an effective equation of state model for the interstellar medium (ISM) that provides a floor to the sound speed (Booth & Schaye 2009). This leads to a growth history that is too delayed to explain high-redshift luminous AGNs, unless massive seeds are invoked. While IMBHs can still grow significantly via mergers in permissive seeding models that predict frequent seed formation (Bhowmick et al. 2024a), an alternative solution to this problem cold be that the Bondi-Hoyle formula is simply inadequate for this early growth regime. Since neither our theoretical understanding, nor our observational capabilities are able to rule out either scenario directly, we need to model both in cosmological simulations to explore indirect effects.
@@
This work is part of the ‘Learning the Universe’ collaboration111https://www.learning-the-universe.org dedicated to understanding the extragalactic Universe by jointly inferring the initial conditions and physical laws that govern its subsequent evolution. As part of the effort to build the predictive galaxy formation model needed for this endeavour, we present a new, flexible parametrisation for gas inflows and MBH growth via accretion in large-volume cosmological simulations. In the first step, the gas properties are measured at a fixed, galactic scale, assumes a scaling of the mass inflow rate with radius, and feeds a simple prescription for accretion disk evolution, which ultimately grows the MBH. The paper is structured as follows. Section 2 describes our numerical setup and models, Section 3 presents the results. We discuss our findings in Section 4 and present our conclusions in Section 5.
@@ 2.1 Initial conditions
We ran uniform resolution cosmological simulations starting at , created by adopting the Planck Collaboration et al. (2016) cold dark matter cosmological parameters (, , , km s-1 Mpc km s-1 Mpc-1, and ). We also used the same linear theory power spectrum as the IllustrisTNG simulations. The initial displacement is calculated using the NgenIC code in the version originally developed for Gadget4 (Springel et al. 2021) using second-order Lagrangian perturbation theory and applying a variance suppression technique (Angulo & Pontzen 2016) to enhance the representativeness of the simulated volume. The gas was initialised from the same density and velocity field, splitting the matter into an equal number of gas cells and dark matter particles and distributing the mass according to the cosmic baryon and dark matter fractions, respectively. A uniform magnetic field was initialised with co-moving Gauss.
In the following, we present cosmological simulations with two different volumes. Specifially, to understand the overall evolution over cosmic time, we evolved boxes of sidelength Mpc (L50) to a redshift of We probed the structure formation of relatively massive halos, but at comparably low mass resolution. To study the high-redshift, early growth of IMBHs, we evolved smaller volumes with sidelength of Mpc (L12.5) to redshift . These simulations adopted lower MBH seed masses and were used to study the IMBH growth at high redshifts. All simulations were run with , and initial resolution elements, covering a dynamic range of in mass resolution overall and an identical mass resolution of the lowest resolution (L12.5) and the highest resolution (L50) simulation. Unless stated explicitly, we show the highest resolution version of the respective simulation box.
@@ 2.2 Reference: The IllustrisTNG model
The simulations were carried out using the Arepo code (Springel 2010; Pakmor et al. 2011, 2016; Weinberger et al. 2020) solving the magneto-hydrodynamics (MHD) equations using the finite-volume technique on a moving mesh and the Poisson equation using a tree-particle-mesh method. An eight-wave Powell cleaning scheme was applied to ensure approximately divergence-free magnetic fields (Pakmor & Springel 2013).
@@
Our main aim is to develop and test a new model for accretion onto MBHs in cosmological environments. We used the IllustrisTNG galaxy formation model as a well-studied framework with comprehensive analysis of the behaviour of its simulated MBH population in the literature (e.g. Bhowmick et al. 2020; Terrazas et al. 2020; Li et al. 2020; Truong et al. 2021; Habouzit et al. 2021, 2022a, 2022b, among many others). In the following, we briefly summarise the IllustrisTNG model. For a more detailed description, we refer to the IllustrisTNG method papers (Weinberger et al. 2017; Pillepich et al. 2018).
@@
We included radiative cooling down to K using primordial (Katz et al. 1996) and metal-line cooling (Wiersma et al. 2009), taking into account a time dependent spatially homogeneous UV background (Faucher-Giguère et al. 2009)333Note that revised backgrounds exist (Puchwein et al. 2019; Faucher-Giguère 2020). However, to be consistent with IllustrisTNG we use the original model. and adding corrections for self-shielding (Rahmati et al. 2013). Additionally, the radiation field of nearby AGN is taken into account. The implementation details are described in Vogelsberger et al. (2013).
@@
The star formation was modelled following Springel & Hernquist (2003); Vogelsberger et al. (2013), with dense gas above a threshold density of cm-3 established as the star-forming limit. We used an effective equation of state (or in a hotter state), which allowed us to model the unresolved ISM. Star particles form in a stochastic manner from star forming gas and are evolved assuming a single stellar population with Chabrier (2003) initial mass function, taking into account chemical enrichment and mass return from asymptotic giant branch stars, core-collapse and type Ia supernovae (with the TNG model updates as described in Pillepich et al. 2018). Stellar feedback is parametrised via massive, enriched galactic wind particles that are temporarily decoupled from the gas dynamics (Springel & Hernquist 2003; Vogelsberger et al. 2013) using the scaling of the IllustrisTNG simulation (Pillepich et al. 2018).
@@
The MBH modelling follows a relatively simple, numerically robust approach for seeding, dynamics, and mergers (Springel et al. 2005; Sijacki et al. 2007; Di Matteo et al. 2008). Specifically, in every halo with a mass exceeding M⊙ that does not yet contain one, a MBH is seeded converting the most bound gas cell into the BH particle with seed mass M⊙. This BH particle is kept at the potential minimum of the halo throughout the simulation, with MBH mergers being modelled instantaneously without explicitly following the dynamics of BHs.
@@
In the TNG model, gas accretion onto MBHs is modelled using the Eddington-limited Bondi formula applied to a kernel-averaged estimate of the gas density, , and sound speed, , in the cells surrounding the BH particle (average denoted by ). This equations are expressed as
@@
where is the gravitational constant, the MBH mass, the proton mass, the radiative efficiency, the Thompson cross-section and the speed of light. The corresponding BH mass growth is
@@
The averaging scale, is adjusted on the fly at every time step using a target weighted number of cells, as
@@
with a given tolerance444Note that we keep this number fixed with resolution, which is a deviation from the IllustrisTNG simulation, where it is scaled by a factor of for each resolution step (factor of in target mass). and where denotes the distance of the cell mesh-generating point and the MBH position and the target gas mass (i.e. the gas mass resolution). The weighting kernel is defined as
@@
Feedback from AGNs is modelled in two modes (Weinberger et al. 2017). Highly accreting MBHs relative to the Eddington limit inject thermal energy continuously into their surroundings (using the same kernel)
For low-accretion-rate MBHs, a kinetic kick is injected in a pulsed fashion whenever enough energy is available. This is expressed as
We note that both feedback channels are mutually exclusive, with an Eddington ratio threshold of